Toombs Media Corp. recently completed a 3-for-1 stock split. Prior to the split, its stock sold for $120 per share. The firm's total market value was unchanged by the split. Other things held constant, what is the best estimate of the stock's post-split price? 


a. $20.00 b. $25.00 c. $30.00 d. $38.73 e. $40.00
Answer #1

3 for 1 split implies that , one will get 3 stocks for each stock

let there be 100 stocks before split

=>

Old value = 100 * 120

New value = 100 * 3 * x

=>

new price 'x' = 100*120/100*3

= 120/3

= 40

hence choose e)